OpenStudy (anonymous):

Find the solutions to the nonlinear equations y=4x x^2+y^2=68

OpenStudy (anonymous):

substitute y=4x to x^2+y^2=68 x^2+(4x)^2=68 x^2 + 16x^2 = 68 17x^2 = 68 x^2 = 4 x=+2 or -2